For prolonged surgical procedures, succinylcholine has been given by IV infusion in order to sustain muscle relaxation. The usual initial dose is 20 mg followed by continuous infusion of 0.5 to 5.0 mg/min. The infusion must be individualized because of great variation in the metabolism (a bipolar distribution of slow and fast metabolizers exists.) Estimate the elimination half-lives of succinylcholine in the patients requiring 0.5 mg/min and 5.0 mg/min infusions to maintain 20 mg of drug in the body.
Answer #1

20 mg of drug has to be maintained in the body.

At 0.5 mg/ min

T1/2 = 20/0.5 = 40 minutes

At 5 mg/min

T1/2 = 20/5 = 4 minutes